(36)Which of the following characteristics about command economies is false?

(a)There is voluntary exchange between buyers and sellers in markets

(b)Profit is the main motivation for economic activity

(c)Social ownership is the dominant form of property ownership

(d)Central planning is the dominant economic system

(37)Suppose the demand and supply equations in the market for fresh milk are given as D=340-3P and S=260+7P in the City of Los Angeles, what is the market equilibrium price and the market equilibrium quantity for fresh milk in this market?

(a)P=$6 ,Q= 412 liters

(b)P=$12,Q=520 liters

(c)P=$8,Q=316 liters

(d)P=$9,Q=285 liters

(38)Suppose now the City government in Los Angeles intervenes in the market for fresh milk referenced in Q#37 above and sets a price of $5 for a gallon of fresh milk, the pricing policy of the City government is called:

(a)A binding price floor

(b)A non-binding price ceiling

(c)A non-binding price floor

(d)A do nothing price

(39)Which of the following statements about the pricing policy adopted by the City government of Los Angeles in the market for fresh milk cited in Q#38 above is false?

(a)The pricing policy is set below the competitive equilibrium price of fresh milk

(b)The pricing policy creates more demand and less supply of fresh milk

(c)The pricing policy is set above the competitive equilibrium price for fresh milk

(d)The pricing policy will not hinder the capacity of the market forces from clearing the market
Answer #1

a) "D"

Central planing are the dominant economic system in the command economy. and everything is owned by the government.

b) "C"

The price in the market at equilibrium is 8 and the quantity demand is 316 liters. we can get the answer by making an equation were the demand and the supply are equal .that is 340 -3p = 260+7p = 340-260 = 10P = p = 8. hence the answer is C.

c) As the price is below the equilibrium price it will be considered as binding price ceiling. answer is "B" a binding price ceiling (i think the option given there is wrong.)

d) "C"

Statement C that the price is set above the competitive price is false.